### v1

# Procountable groups are not classifiable by countable structures

Abstract.
We prove that the relation of topological isomorphism on procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ that expresses that two sequences of reals have bounded difference is Borel reducible to it. This marks progress on an open problem of [15], to determine the exact complexity of isomorphism among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-1931. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this work was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (Consolidator grant) “Groups, Crystals and Classifications”.
1. Introduction
Suppose we are given a concrete presentation of abstract mathematical objects of a certain type. It is of interest to know the complexity of determining whether two presentations describe the same abstract object. For instance, finite graphs can be concretely described by binary strings that encode adjacency matrices. The graph isomorphism problem asks whether two such strings describe the same graph. This problem is in NP, and its exact complexity has been the object of much recent study [1, 30]. The triviality problem for groups is to determine whether a finite presentation of a group decribes the trivial group. This is undecidable [29]. Therefore the $\Sigma^{0}_{1}$ relation of isomorphism between finitely presented groups is undecidable. In fact every $\Sigma^{0}_{1}$ equivalence relation is computably reducible to it ([22]; also see [25, Cor. 2.6]).
The presentations in the previous two examples are finite. However, for many classes of mathematical objects, a countably infinite amount of information is required to describe a member. In this case, one can usually view the descriptions as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in [7], is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation [8] denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
The programme of [15]. A Polish group is a topological group whose topology is Polish, i.e., separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups [2]. Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ [2]. It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space (see [15]). So we can take the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ as the concrete presentations of non-archimedean Polish groups. In Kechris et al. [15] the following programme was initiated:
Given a Borel class $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$, determine the complexity of topological isomorphism between groups in $\mathcal{C}$.
There has been some progress on this program. We summarize the results that are known to us below.
  • One class considered in [15] was the compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bireducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel.
  • One can also consider the subclass of abelian profinite groups. Due to the Pontryagin duality, the isomorphism relation of abelian profinite groups is Borel bireducible with the isomorphism of countable torsion abelian groups. It is well known that the latter relation is classified by the Ulm invariants, and is known to be a complete analytic set and hence non-Borel (see [7, (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$.
  • The isomorphism relation of oligomorphic groups was shown to be essentially countable in [23]; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see [24, 26]; in particular in [26] it was shown that under the assumption of weak elimination of imaginaries the relation is smooth.
  • The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in [4]. Work in progress of Ruiwen Li would show that this relation is Borel bireducible with GI.
Our work marks considerable progress on an open problem of Kechris et al. [15] pertaining to this programme. In [15], the authors gave a criterion determining when isomorphism on an invariant Borel class of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ can be classified by countable structures. They asked at the end of their introduction for the complexity of isomorphism for the class of all non-archimedean Polish groups, observing that this relation is analytic and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of procountable groups already has a complexity much higher than GI.
We note that the programme of [15] also asked to first determine whether a class $\mathcal{C}$ of interest is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. The Borelness is unknown for the class of topologically finitely generated groups, and known to be false for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be coanalytic and non-Borel in [18], and in [28] it is shown that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete coanalytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
The main result.
Definition 1.1.
A topological group $G$ is called procountable if $G$ is topologically isomorphic to the inverse limit of an inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ of countable groups and onto homomorphisms $p_{n}\colon G_{n+1}\to G_{n}$, where the topology of $G$ is the subspace topology given by the product $\prod_{n}G_{n}$, where each $G_{n}$ is equipped with the discrete topology.
It is easily seen that each procountable group is non-archimedean Polish, thus it is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. We will verify in 2.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of $1$ of normal open subgroups (see, e.g., [12, Theorem 3.3] or [19, Lemma 2]). Since each countable discrete group is procountable, we know that GI is Borel reducible to the isomorphism of procountable groups through any of the following results about countable groups:
  • The isomorphism of countable locally finite groups that are nilpotent of class 22 is Borel bireducible with GI (Mekler [21], who proves this in fact for groups of exponent 3; also see [7, (2.3)]);
  • The isomorphism of countable torsion-free abelian groups is Borel bireducible with GI ([27]);
  • The isomorphism of countable omnigenous locally finite groups is Borel bireducible with GI ([10]).
In particular, we get that the isomorphism relation of procountable groups is a complete analytic set and hence non-Borel. In this paper we address the problem of its complexity in the Borel reducibility hierarchy.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality between sequences of real numbers. By a celebrated theorem of Kechris and Louveau [14], $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation which Borel reduces $E_{1}$ has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
$(x_{n})\ell_{\infty}(y_{n})\iff\exists M\in\mathbb{R}\ \forall n\in\mathbb{N}\ |x_{n}-y_{n}|<M.$
By a theorem of Rosendal (see e.g. [8, Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
Theorem 1.2.
The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on procountable groups. Consequently, the topological isomorphism relation on procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
It is known that being Borel reducible to GI is equivalent to being below the orbit equivalence relation of a Borel action of $\mathrm{Sym}(\mathbb{N})$, which is also equivalent to being classifiable by some sort of countable structures. It follows from our main theorem that the topological isomorphism of procountable groups is not classifiable by countable structures.
The proof makes use of the notion of uniform homeomorphism.
Definition 1.3.
A Polish metric space is a metric space $(X,d)$ where $d$ is a complete metric and the topology generated by $d$ is Polish. Let $X$ and $Y$ be Polish metric spaces. A bijection $\Phi$ between $X$ and $Y$ is called a uniform homeomorphism if both $\Phi$ and its inverse are uniformly continuous. $X$ and $Y$ are said to be uniformly homeomorphic if there is a uniform homeomorphism between $X$ and $Y$.
The two steps of the proof of the main result 1.2.
  1. 1.
    Section 3 shows that $\ell_{\infty}$ is Borel reducible to the relation of uniform homeomorphism between path spaces $[T]$, where $T$ is a pruned subtree of $\mathbb{N}^{<\mathbb{N}}$ and $[T]$ carries the standard ultrametric.
  2. 2.
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the space of procountable groups.
The second step only uses a very particular type of groups that could be called “solenoid groups”. Any group $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ is isomorphic to a free Coxeter group generated by an infinite set $V_{n}$, namely $\langle v\in V_{n}\mid v^{2}=e\rangle$. The binding maps are given by surjections $V_{n+1}\to V_{n}$.
Open questions.
We close the introduction with the following open problems. The first two will be discussed again in see Section 3.4.
Question 1.4.
Is the relation of topological isomorphism on procountable groups a universal analytic equivalence relation (i.e., any analytic equivalence relation is Borel reducible to it)?
It is known that the relation of uniform homeomorphism on Polish metric spaces is universal analytic [5] .
Question 1.5.
Is the relation of uniform homeomorphism on ultrametric Polish spaces a universal analytic equivalence relation?
By the second step above, an affirmative answer to the second question would also answer the first in the affirmative.
Question 1.6.
Is there a nilpotent variety of groups $\mathbf{V}$ such that the relation of topological isomorphism on procountable groups in $\mathbf{V}$ is not classifiable by countable structures? In particular, are abelian procountable groups classifiable by countable structures?
It follows from [27] that GI is a lower bound for topological isomorphism on procountable abelian groups.
2. Presentations of procountable groups
Recall Definition 1.1. Encoding inverse systems by countable structures with infinitely many sorts, we obtain a standard Borel space $\mathcal{P}$ of procountable groups. We show in this section that these presentations can be converted in a Borel way into procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$, and that the latter form a Borel set $\mathcal{C}$.
Notation 2.1.
The space of all finite sequences of natural numbers is denoted as $\mathbb{N}^{<\mathbb{N}}$. For $s,t\in\mathbb{N}^{<\mathbb{N}}$ and $x\in\mathbb{N}^{\mathbb{N}}$, we use $s\subseteq t$ to denote that $s$ is an initial segment of $t$ and similarly for $s\subseteq x$. For $s\in\mathbb{N}^{<\mathbb{N}}$, we write
$s]=\left\{x\in\mathrm{Sym}(\mathbb{N})\colon\,s\subseteq x\right\}.$
For a closed subgroup $G\leqslant_{c}\mathrm{Sym}(\mathbb{N})$, we will write $s]_{G}$ for $s]\cap G$. We let $\sigma_{n}$ denote the tuple $(0,1,\ldots,n-1)$; note that the $\sigma_{n}]_{G}$ form a neighbourhood basis of the identity in $G$ consisting of open subgroups.
Given $s,t\in\mathbb{N}^{<\mathbb{N}}$, let $s\circ t$ be the sequence representing composition, on the longest initial segment of $\mathbb{N}$ in which it is defined. Let $s^{-1}$ be the inverse of $s$ on the longest initial segment in which it is defined. For instance, $(7,4,3,1,0)\circ(3,4,6)=(1,0)$, and $(1,2,0,5)^{-1}=(2,0,1)$.
[…]
$G=\left\{x\in\prod_{n}G_{n}\colon\,\forall n\,[p_{n}(x(n+1))=x(n)]\right\}.$
Remark 2.2.
From an inverse system as above, one can in a Borel way determine a corresponding closed subgroup of $\mathrm{Sym}(\mathbb{N})$. One replaces $G_{n}$ by an isomorphic group with domain $D_{n}=\{\langle i,n\rangle\colon i\in\mathbb{N}\}$, where $\langle\cdot,\cdot\rangle\colon\mathbb{N}\times\mathbb{N}\to$$\mathbb{N}$ is the Cantor pairing function. Then the left regular presentation of $G_{n}$ is a copy of $G_{n}$ as a permutation group with support $D_{n}$, and the product of these copies in $\mathrm{Sym}(\mathbb{N})$ is a copy $C$ of $\prod_{n}G_{n}$ as a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. Finally, one obtains a copy of $G$ as a closed subgroup of $C$ in a Borel way from the binding maps $p_{n}$.
Proposition 2.3.
The procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel class $\mathcal{C}$.
Proof.
As mentioned before, $G$ is procountable if and only if the identity has a neighbourhood basis of normal open subgroups. In the setting of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ this is equivalent to: for each $n$, there is $k>n$ such that the normal closure (in $G$) of $\sigma_{k}]_{G}$ is contained in $\sigma_{n}]_{G}$. Since $\sigma_{n}]_{G}$ is a subgroup, it suffices to express as a Borel condition about $G$ that
(1) $\forall x\in G\,\left[x^{-1}\circ([\sigma_{k}]_{G})\circ x\subseteq[\sigma_{n}]_{G}\right].$
In the following, letters $s,t$ etc. range over sequences that have an extension in $G$. Recalling 2.1, we claim that (1) is equivalent to the Borel condition
(2) $\forall s\ \forall t\supseteq\sigma_{k}\ \ [s^{-1}\circ t\circ s]_{G}\cap[\sigma_{n}]_{G}\neq\emptyset.$
For (1) $\to$ (2), given $s$ and $t\supseteq\sigma_{k}$, let $x\supseteq s$, $x\in G$ and $y\supseteq t$, $y\in G$. Then $x^{-1}\circ y\circ x\in[\sigma_{n}]_{G}$.
For (2) $\to$ (1), suppose that there is $x\in G$ such that $x^{-1}\circ([\sigma_{k}]_{G})\circ x$ is not contained in $\sigma_{n}]_{G}$. Then, by the continuity of the conjugation operation, for sufficiently long $s\subseteq x$ and sufficiently long $t\supseteq\sigma_{k}$, we have $s^{-1}\circ t\circ s]_{G}\cap[\sigma_{n}]_{G}=\emptyset$. ∎
Remark 2.4.
[…]
3. Borel reduction of $\ell_{\infty}$ to uniform homeomorphism
Definition 3.1.
By a tree on $\mathbb{N}$ we mean a nonempty set $T\subseteq\mathbb{N}^{<\mathbb{N}}$ that is closed under initial segments. The path space
$T]=\big\{x\in\mathbb{N}^{\mathbb{N}}\colon\forall n\in\mathbb{N}\ (x\!\!\upharpoonright\!n\in T)\big\}$
is a Polish ultrametric space with the metric given by declaring for $x\neq y$ that
[…]
The elements of $[T]$ will be called branches of $T$.
A tree $T$ is called pruned if for any $s\in T$ there is $t\in T$ such that $s\subsetneq t$. The space of all pruned trees on $\mathbb{N}$ can be seen to be a standard Borel (even Polish) subspace of $2^{(\mathbb{N}^{<\mathbb{N}})}$, which will be denoted $\mathcal{T}$.
Note that the paths ultrametric spaces are in a canonical bijection with the pruned trees, so we can see the elements of $\mathcal{T}$ as codes for such spaces.
Definition 3.2.
By $\cong_{u}$ we denote the equivalence relation of uniform homeomorphism on $\mathcal{T}$.
The following quick observation gives a lower bound on the complexity of $\cong_{u}$.
Proposition 3.3.
GI is Borel reducible to $\cong_{u}$.
Proof.
We use the main result of [3], where it was shown that GI is Borel bireducible with the homeomorphism relation between 0-dimensional compact metric spaces. Let $\Phi$ denote a Borel witness to this reduction. Then for any graphs $\Gamma,\Gamma^{\prime}$ with vertex set $\mathbb{N}$, $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $\Phi(\Gamma)$ and $\Phi(\Gamma^{\prime})$ are homeomorphic. Note that a Polish space is 0-dimensional if and only if it admits a compatible ultrametric. So there exists a compatible ultrametric $d_{\Gamma}$ for $\Phi(\Gamma)$. By the standard tools from descriptive set theory, we may find a Borel map giving the correspondence $\Phi(\Gamma)\mapsto d_{\Gamma}$.
Note that for any compact ultrametric space $(X,d_{X})$, the set of nonzero distances $\{d_{X}(x,y)\colon x,y\in X\}\setminus\{0\}$ is a decreasing sequence of positive real numbers converging to 0. Up to homeomorphism, we may assume that this sequence is $\{2^{-n}\colon n\in\mathbb{N}\}$. Now that $(\Phi(\Gamma),d_{\Gamma})$ is a compact ultrametric space with its set of nonzero distances a subset of $\{2^{-n}\colon n\in\mathbb{N}\}$, there is a unique tree $T_{\Gamma}\in\mathcal{T}$ such that $(\Phi(\Gamma),d_{\Gamma})$ is isometric to $T_{\Gamma}]$. Again the map $d_{\Gamma}\mapsto T_{\Gamma}$ is Borel.
Finally, note that, as compact spaces, $T_{\Gamma}]$ and $T_{\Gamma^{\prime}}]$ are homeomorphic if and only if they are uniformly homeomorphic. Thus we have that $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $T_{\Gamma}]$ and $T_{\Gamma^{\prime}}]$ are uniformly homeomorphic as compact ultrametric spaces. This proves the proposition. ∎
It follows that $\cong_{u}$ is a complete analytic set and hence non-Borel.
The rest of this section will establish the following result of independent interest, which is the first step towards proving our Main Result 1.2.
Theorem 3.4.
$\ell_{\infty}$ is Borel reducible to $\cong_{u}$, namely, uniform homeomorphism of the ultrametric spaces denoted by points in $\mathcal{T}$. Consequently, $\cong_{u}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, $\cong_{u}$ is not Borel reducible to GI.
3.1. Some facts on uniform homeomorphisms
The following is a direct consequence of the definition of uniform continuity.
Lemma 3.5.
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be continuous. Then $\Phi$ is uniformly continuous if and only if for all $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that for any $x,y\in[T]$, if $x\!\!\upharpoonright\!m=y\!\!\upharpoonright\!m$, then $\Phi(x)\!\!\upharpoonright\!n=\Phi(y)\!\!\upharpoonright\!n$.
In this context we define $\Omega_{\Phi}(n)$ to be the least $m$ such that the condition in the above lemma holds. Note that the lemma implies
$\big|\big\{t\in T\colon|t|=\Omega_{\Phi}(n)\big\}\big|\geqslant\big|\big\{s\in S\colon|s|=n\big\}\big|.$
Lemma 3.6.
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be a uniform homeomorphism (see 1.3). If both $[T],[S]$ are uncountable, then $\Omega_{\Phi}$ and $\Omega_{\Phi^{-1}}$ are unbounded monotone functions from $\mathbb{N}$ to $\mathbb{N}$.
Proof.
The monotonicity is obvious. If $\Omega_{\Phi}$ is bounded by $m$, then for each $n$ the string $\Phi(x)\!\!\upharpoonright\!n$ is determined by $x\!\!\upharpoonright\!m$. So $[S]$ is countable. ∎
3.2. Definition of the reduction
We will need to restrict the entries of the sequences in $\mathbb{R}^{\mathbb{N}}$ to natural numbers. This does not decrease the complexity of the equivalence relation:
Lemma 3.7.
$\ell_{\infty}\leqslant_{B}\ell_{\infty}\!\!\upharpoonright\!\mathbb{N}^{\mathbb{N}}$.
Proof.
Define $f\colon\mathbb{R}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$ by
$f(x)(2n)=\left\{\begin{array}[]{ll}\lfloor x(n)\rfloor,&\mbox{if $x(n)\geqslant 0$,}\\ 0,&\mbox{otherwise,}\end{array}\right$.
and
$f(x)(2n+1)=\left\{\begin{array}[]{ll}0,&\mbox{if $x(n)\geqslant 0$,}\\ -\lfloor x(n)\rfloor,&\mbox{otherwise.}\end{array}\right$.
Then for any $x,y\in\mathbb{R}^{\mathbb{N}}$ and $n\in\mathbb{N}$, we have $|f(x)(n)-f(y)(n)|\leqslant|x(n)-y(n)|$. Thus, if $x\ell_{\infty}y$, then $f(x)\ell_{\infty}f(y)$. Conversely, if |f(x)(n)f(y)(n)|<M|f(x)(n)-f(y)(n)|<M for all $n\in\mathbb{N}$, then for all $n\in\mathbb{N}$, either $x(n)y(n)>0$ and |x(n)y(n)|<M|x(n)-y(n)|<M hold, or else $x(n)y(n)\leqslant 0$ and $|x(n)|,|y(n)|<M$ hold; in both cases we have |x(n)y(n)|<2M|x(n)-y(n)|<2M. Thus $f$ is a reduction from $\ell_{\infty}$ to $\ell_{\infty}\!\!\upharpoonright\!\mathbb{N}^{\mathbb{N}}$. ∎
For the next step of the reduction, we consider an arbitrary $x\in\mathbb{N}^{\mathbb{N}}$ and define a pruned tree $\mathsf{T}_{x}$ on $\mathbb{N}$. Before defining $\mathsf{T}_{x}$, we first introduce some auxiliary trees
$\mathsf{S}_{*},\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots,\mathsf{S}_{\omega},\mathsf{R}_{0},\mathsf{R}_{1},\dots,\mathsf{R}_{k},\dots$
that will be used as building blocks. Let $\mathsf{S}_{*}\subseteq\mathbb{N}^{<\mathbb{N}}$ be the tree consisting of all finite sequences $s\in\mathbb{N}^{<\mathbb{N}}$ such that $s(n)$ is even for each $n\in\mathbb{N}$. Then $\mathsf{S}_{*}$ is a copy of the full tree $\mathbb{N}^{<\mathbb{N}}$. The point of defining $\mathsf{S}_{*}$ is that we may arbitrarily extend $\mathsf{S}_{*}$ by adding extensions of existing nodes, and still maintain that the extended tree be an “official” subtree of $\mathbb{N}^{<\mathbb{N}}$.
The trees $\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots$ will be abstractly and inductively defined. Their exact realizations as subtrees of $\mathbb{N}^{<\mathbb{N}}$ will not be important. To begin, let $\mathsf{S}_{1}$ be a single branch. Thus the Cantor–Bendixson rank of $\mathsf{S}_{1}]$ is $1$. Let $\mathsf{S}_{2}$ be a pruned tree so that the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $\mathsf{S}_{1}]$. So the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $2$. In general, let $\mathsf{S}_{k+1}$ be a pruned tree whose Cantor–Bendixson derivative is $\mathsf{S}_{k}]$. Then the Cantor–Bendixson rank of $\mathsf{S}_{k+1}$ is $k+1$. This finishes the definitions of $\mathsf{S}_{k}$ for $1\leqslant k<\mathbb{N}$.
Next we define $\mathsf{S}_{\omega}$ to be the extension of $\mathsf{S}_{*}$ defined as follows. For each $k\in\mathbb{N}$, let $\mathsf{z}_{k}$ be the branch of $\mathsf{S}_{*}$ defined as
$\mathsf{z}_{k}(m)=\left\{\begin{array}[]{ll}2k,&\mbox{if $m=0$,}\\ 0,&\mbox{otherwise.}\end{array}\right$.
For any $k,m\in\mathbb{N}$, let $\mathsf{t}_{k,m}=\mathsf{z}_{k}\!\!\upharpoonright\!(m+1)$. $\mathsf{S}_{\omega}$ is obtained by extending $\mathsf{S}_{*}$ with appending a copy of $\mathsf{S}_{k}$ at the node $\mathsf{t}_{k,2m}$ for all $k,m\in\mathbb{N}$.
The distinctive feature of $\mathsf{S}_{\omega}$ is that any autohomeomorphism of $\mathsf{S}_{\omega}]$ must keep $\mathsf{z}_{k}$ fixed for all $k\in\mathbb{N}$. This is because of the topological characterization of $\mathsf{z}_{k}$ as the only $z\in[\mathsf{S}_{\omega}]$ with the property that:
for any clopen set $U$ with $z\in U$, the $k$-th Cantor–Bendixson derivative of $U$ is the perfect kernel of $U$, and for any $i<k$, the $i$-th Cantor–Bendixson derivative of $U$ contains a perfect subset but is not itself perfect.
Next we abstractly define a sequence of finitely splitting pruned trees
$\mathsf{R}_{0},\mathsf{R}_{1},\dots,\mathsf{R}_{k},\dots$
so that no $\mathsf{R}_{k}]$ has an isolated point, as follows. For each $k\in\mathbb{N}$, let $\mathsf{R}_{k}$ be $2^{k}$-splitting on the first level, and then each node on the first level be extended with a copy of the full binary tree. It is easy to see that for any $k\in\mathbb{N}$, $\mathsf{R}_{k}]$ is homeomorphic to a Cantor set, and therefore there is a uniform homeomorphism between $\mathsf{R}_{k}]$ and $\mathsf{R}_{\ell}]$ for any $k,\ell\in\mathbb{N}$.
Finally we are ready to define $\mathsf{T}_{x}$ given an $x\in\mathbb{N}^{\mathbb{N}}$. $\mathsf{T}_{x}$ is obtained from $\mathsf{S}_{\omega}$ by extending the node $\mathsf{t}_{k,2m+1}$ by a copy of $\mathsf{R}_{x(k)}$ for all $k,m\in\mathbb{N}$. This finishes the definition of $\mathsf{T}_{x}$.
For each $x\in\mathbb{N}^{\mathbb{N}}$, we let $\mathsf{S}_{*}(\mathsf{T}_{x})$, $\mathsf{S}_{\omega}(\mathsf{T}_{x})$, $\mathsf{t}_{k,m}(\mathsf{T}_{x})$, $\mathsf{z}_{k}(\mathsf{T}_{x})$ etc. denote respectively the subtrees $\mathsf{S}_{*}$, $\mathsf{S}_{\omega}$, nodes $\mathsf{t}_{k,m}$, and branches $\mathsf{z}_{k}$ etc. in the construction of $\mathsf{T}_{x}$. Note that for different $x$ these objects are isomorphic, but we need to address the specific copies of them as substructures of $\mathsf{T}_{x}$.
For $k,m\in\mathbb{N}$, we let $\mathsf{R}_{x(k),m}$ denote the copy of $\mathsf{R}_{x(k)}$ which was appended at the node $\mathsf{t}_{k,2m+1}(\mathsf{T}_{x})$.
Note that for any $x\in\mathbb{N}^{\mathbb{N}}$ and $k\in\mathbb{N}$, $\mathsf{z}_{k}(\mathsf{T}_{x})$ is still topologically characterized by the same property as above, since the additional sets added beyond $\mathsf{S}_{\omega}]$ in the construction of $\mathsf{T}_{x}]$ are all perfect.
Therefore, if $\theta$ is any homeomorphism from $\mathsf{T}_{x}]$ to $\mathsf{T}_{y}]$, then we must have that $\theta(\mathsf{z}_{k}(\mathsf{T}_{x}))=\mathsf{z}_{k}(\mathsf{T}_{y})$ for all $k\in\mathbb{N}$.
Although we do not need this in the proof, we observe that for any $x,y\in\mathbb{N}^{\mathbb{N}}$, $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$ are homeomorphic.
Definition 3.8.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be Polish metric spaces and let $f\colon X\to Y$. Let $L>0$. We say that $f$ is $L$-Lipschitz if for any $x,x^{\prime}\in X$,
$d_{Y}(f(x),f(x^{\prime}))\leqslant L\,d_{X}(x,x^{\prime}).$
$f$ is Lipschitz if it is $L$-Lipschitz for some $L>0$.
For $k,\ell\in\mathbb{N}$, consider the natural homeomorphism $\Phi_{k,\ell}$ between $\mathsf{R}_{k}]$ and $\mathsf{R}_{\ell}]$. For later use we note that $\Phi_{k,\ell}$ is $L$-Lipschitz with $L=2^{\ell-k}$. Similarly, $\Phi_{k,\ell}^{-1}=\Phi_{\ell,k}$ is $L^{-1}$-Lipschitz.
3.3. Why the reduction works
Assume $x,y\in\mathbb{N}^{\mathbb{N}}$ with $x\ell_{\infty}y$. Let $M>0$ be such that |x(n)y(n)|<M|x(n)-y(n)|<M for all $n\in\mathbb{N}$. We define a uniform homeomorphism $\Phi$ between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$. For this, note that $\mathsf{T}_{x}]$ is the disjoint union of $\mathsf{S}_{\omega}(\mathsf{T}_{x})]$ and $\mathsf{R}_{x(k),m}]$ for all $m\in\mathbb{N}$, and $\mathsf{T}_{y}]$ is the disjoint union of $\mathsf{S}_{\omega}(\mathsf{T}_{y})]$ and $\mathsf{R}_{y(k),m}]$ for all $m\in\mathbb{N}$. Let $\Phi$ be the homeomorphism from $\mathsf{S}_{\omega}(\mathsf{T}_{x})]$ onto $\mathsf{S}_{\omega}(\mathsf{T}_{y})]$ induced by the identity map from $\mathsf{S}_{\omega}(\mathsf{T}_{x})$ to $\mathsf{S}_{\omega}(\mathsf{T}_{y})$. Then it is obvious that $\Phi$ is an isometry, and therefore both $\Phi$ and $\Phi^{-1}$ are $1$-Lipschitz. Let $\psi_{k,m}$ be the natural homeomorphism between $\mathsf{R}_{x(k),m}]$ and $\mathsf{R}_{y(k),m}]$. Then for all $k,m\in\mathbb{N}$, both $\psi_{k,m}$ and $\psi_{k,m}^{-1}$ are $2^{M}$-Lipschitz. Let $\Psi=\Phi\cup\bigcup_{k,m}\psi_{k,m}$. Then both $\Psi$ and $\Psi^{-1}$ are $2^{M}$-Lipschitz. Thus $\Psi$ is a uniform homeomorphism between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$.
Conversely, assume $\Psi$ is a uniform homeomorphism between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$. In particular, $\Psi$ is a homeomorphism. Then $\Psi(\mathsf{z}_{k}(\mathsf{T}_{x}))=\mathsf{z}_{k}(\mathsf{T}_{y})$ for any $k\in\mathbb{N}$. By Lemma 3.6, there is $m>1$ such that
$\Omega_{\Psi}(2m+2)>1$
and, letting $n=\Omega_{\Psi}(2m+2)$, also
$\Omega_{\Psi^{-1}}(2n+2)>1.$
Let $M$ be the least such that $2M+2\geqslant\Omega_{\Psi^{-1}}(2n+2)$. We claim that for any $k\in\mathbb{N}$, $\Psi([\mathsf{R}_{x(k),n}])$ is contained in
$\bigcup_{m<p\leqslant M}[\mathsf{R}_{y(k),p}].$
To see this, let $u\in[\mathsf{R}_{x(k),n}]$ be arbitrary. Then
$u\upharpoonright(2n+1)=\mathsf{z}_{k}(\mathsf{T}_{x})\upharpoonright(2n+1).$
Since $2n+1>n=\Omega_{\Psi}(2m+2)$, we have that
$\Psi(u)\upharpoonright(2m+2)=\mathsf{z}_{k}(\mathsf{T}_{y})\upharpoonright(2m+2).$
However, if $\Psi(u)=v\in[\mathsf{R}_{y(k),p}]$ for $p>M$, then
$v\upharpoonright(2p+1)=\mathsf{z}_{k}(\mathsf{T}_{y})\upharpoonright(2p+1),$
and since $2p+1>2M+2\geqslant\Omega_{\Psi^{-1}}(2n+2)$, we have
$u\upharpoonright(2n+2)=\mathsf{z}_{k}(\mathsf{T}_{x})\upharpoonright(2n+2),$
which is a contradiction.
By our construction, the number of nodes at level $2n+2$ in the copy of $\mathsf{R}_{x(k),n}$ is $2^{x(k)}$. The number of nodes at level $2M+2$ in the copies of $\bigcup_{m<p\leqslant M}\mathsf{R}_{y(k),p}$ is no more than
$2^{y(k)+(M-m)}+2^{y(k)+(M-m-1)}+\cdots+2^{y(k)}=2^{y(k)}(2^{M-m+1}-1)\leqslant 2^{y(k)+M-m+1}.$
By Lemma 3.5 we have that $2^{y(k)+M-m+1}\geqslant 2^{x(k)}$, and hence $x(k)-y(k)\leqslant M-m+1$. This means that x(k)y(k)x(k)-y(k) is bounded. By symmetry, y(k)x(k)y(k)-x(k) is also bounded. Therefore $x\ell_{\infty}y$.
3.4. Further remarks
Definition 3.9.
We call an analytic equivalence relation universal if any analytic equivalence relation is Borel reducible to it.
Such equivalence relations were called complete analytic equivalence relations in [5]. Here we call them universal to avoid possible confusion. In [5] and subsequent research, many natural equivalence relations were found to be universal analytic equivalence relations. In particular, it was shown in [5] that the uniform homeomorphism for all Polish metric spaces is a universal analytic equivalence relation.
Here we remark that our result does not show that $\cong_{u}$ is universal. By 3.3 GI is reducible to $\cong_{u}$, and by the main result 1.2 $\ell_{\infty}$ is reducible to $\cong_{u}$. However, the equivalence relation $c_{0}$ does not reduce to $\mbox{\rm GI}\times\ell_{\infty}$. Recall that $c_{0}$ is the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined as
$(x_{n})c_{0}(y_{n})\iff\lim_{n}|x_{n}-y_{n}|=0.$
Toward a contradiction, assume there was a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [13, Theorem 3.18 and Example 3.23]; also see [8, Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [8, Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [8, Theorem 8.5.2] and its proof).
We conjecture that $\cong_{u}$, and hence uniform homeomorphism of Polish ultrametric spaces, are universal analytic equivalence relations.
4. Borel reduction of uniform homeomorphism to isomorphism of procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, namely we exhibit a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of procountable groups.
For a countably infinite set $V$, we define the free product of cyclic groups of size $2$ with generators in $V$ by
$L(V)=\langle s\mid s^{2}=e\ \ (s\in V)\rangle.$
(This group is also known as the free Coxeter group of rank $\aleph_{0}$.) By the general theory of free products of group (see, e.g., [17, IV.4.7]), each involution is conjugate to a generator, and no two generators are conjugate. If $r\colon V\to W$, then we have a map $\widehat{r}\colon L(V)\to L(W)$ induced by $v\mapsto r(v)$.
Given a pruned tree $T$, let $T_{n}=\{s\in T\colon|s|=n\}$ be the $n$-th level of the tree $T$, and let $r_{n}\colon T_{n+1}\to T_{n}$ be the predecessor map. We may assume that $T_{1}$ is infinite because the trees we obtain through the coding in 3.4 have that property. Let
[…]
We view $G_{T}$ concretely as a closed subgroup of $\mathrm{Sym}(\mathbb{N})$ according to Remark 2.2. Needless to say, the map sending $T$ to $G_{T}$ is Borel. The rest of this section will establish the following, which is the second step towards proving the Main Result 1.2.
Theorem 4.1.
Let $T$ and $U$ be pruned trees on $\mathbb{N}$. Then $T]\cong_{u}[U]$ if and only if $G_{T}$ is topologically isomorphic to $G_{U}$.
4.1. Proω categories
We need some preliminaries on isomorphism of procountable groups corresponding to certain isomorphisms of inverse systems of type $\mathbb{N}$. We mainly give them to provide a notational framework; more general results are obtained in [20].
Definition 4.2.
Let $\mathcal{K}$ be a category that has as objects countable structures in the same signature, and as morphisms surjective homomorphisms between. The category $\mathrm{Pro}_{\omega}(\mathcal{K})$ has as objects the inverse systems $(\overline{A},\overline{p})=(A_{n},p_{n})_{n\in\mathbb{N}}$ over $\mathcal{K}$, where $p_{n}\colon A_{n+1}\to A_{n}$. The pre-morphisms
$\langle\overline{f},\phi\rangle\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$
are given as sequences of morphisms $f_{n}:A_{\phi(n)}\to B_{n}$ where $\phi:\mathbb{N}\rightarrow\mathbb{N}$ is increasing, and all the diagrams are commutative. In more detail, for $k>n$, write $p_{k,n}$ for the “interval composition” maps $p_{n}\circ p_{n+1}\circ\ldots\circ p_{k-1}\colon A_{k}\to A_{n}$ and define $q_{k,n}$ in a similar way; we require that the diagrams
$A_{\phi(k)}$$B_{k}$$A_{\phi(n)}$$B_{n}$$\scriptstyle{f_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{f_{n}}$
are commutative. Composition of pre-morphisms is defined in the obvious way (see [20, pg. 5]) and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=(\mathrm{id}_{A_{n}}:n\in\mathbb{N})$, where of course $\mathrm{id}_{A_{n}}:A_{n}\rightarrow A_{n}$ is the identity map on $A_{n}$.
For each pair of objects $(\overline{A},\overline{p})$ and $(\overline{B},\overline{q})$ we define an equivalence relation $\sim$ between pre-morphisms $(\bar{f},\phi),(\bar{f}^{\prime},\phi^{\prime}):(\overline{A},\overline{p})\to(\overline{B},\overline{q})$. We say that $(\bar{f},\phi)\sim(\bar{f}^{\prime},\phi^{\prime})$ provided for each $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that $m\geqslant\phi(n),\phi^{\prime}(n)$ and the following diagram is commutative:
$A_{m}$$A_{\phi(n)}$$A_{\phi^{\prime}(n)}$$B_{n}$$\scriptstyle{p_{m,\phi(n)}}$$\scriptstyle{p_{m,\phi^{\prime}(n)}}$$\scriptstyle{f_{n}}$$\scriptstyle{f^{\prime}_{n}}$
Transitivity of $\sim$ and compatibility with composition of pre-morphisms is proved in [20, pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are now defined as the equivalence classes of pre-morphisms under $\sim$.
Remark 4.3.
Notice that for $(\bar{f},\phi):(\overline{A},\overline{p})\to(\overline{A},\overline{p})$, we have $(\bar{f},\phi)\sim({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$ if and only if for every $n\in\mathbb{N}$ there is $m\geqslant\phi(n),n$ such that the following diagram commutes:
$A_{m}$$A_{\phi(n)}$$A_{n}$$A_{n}$$\scriptstyle{p_{m,\phi(n)}}$$\scriptstyle{p_{m,n}}$$\scriptstyle{f_{n}}$$\scriptstyle{\mathrm{id}_{A_{n}}}$
that is, $f_{n}:A_{\phi(n)}\rightarrow A_{n}$ is simply the map $p_{\phi(n),n}$.
Let $\mathrm{EpiGrp}$ be the category of countable groups with onto homomorphisms.
Lemma 4.4.
  1. (i)
    The inverse limit operation yields a full functor $P$ from $\mathrm{Pro}_{\mathbb{N}}(\mathrm{EpiGrp})$ to procountable groups with continuous epimorphisms.
  2. (ii)
    $P$ preserves isomorphisms in both directions.
Proof.
[…]
In order to show that the functor $P$ is full, suppose that $\gamma\colon G\to H$ is a continuous epimorphism. We will define a pre-morphism $(\bar{f},\phi)\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$ with $P(\bar{f},\phi)=\gamma$. Let $a_{n}$ and $b_{n}$ be the kernels of the canonical projections $\pi^{G}_{n}:G\to A_{n}$ and $\pi^{H}_{n}:H\to B_{n}$ respectively. Let $\phi$ be an increasing function such that $a_{\phi(n)}\subseteq\gamma^{-1}(b_{n})$. Such $\phi$ exists because $\gamma^{-1}({b_{n}})$ is an open subgroup of $A_{n}$, and the $a_{n}$} form a basis of neighbourhoods of $e$. (This part of the argument is special to procountable groups.) For $n\in\mathbb{N}$ we define $f_{n}:A_{\phi(n)}\rightarrow B_{n}$ via the diagram
[…]
We have to verify commutativity of the following diagram, where $k>n$:
$A_{\phi(k)}$$B_{k}$$A_{\phi(n)}$$B_{n}\ .$$\scriptstyle{f_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{f_{n}}$
In order to do so it suffices to look at the diagram
[…]
and recognize its commutativity.
(ii). For the forward direction, consider an isomorphism
$(\bar{f},\phi)\colon\overline{A}\rightleftarrows\overline{B}:(\bar{g},\psi)$
in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall to equal
$A_{\phi(\psi(n))}\xrightarrow{f_{\psi(n)}}B_{\psi(n)}\xrightarrow{g_{n}}A_{n},$
is such that $(\bar{g},\psi)\circ(\bar{f},\phi)\sim(\mathrm{id}_{\overline{A}},\mathrm{id}_{\mathbb{N}})$. By 4.3, this simply means that $g_{n}\circ f_{\psi(n)}=p_{\phi(\psi(n)),n}$. We argue similarly for the composition $(\bar{f},\phi)\circ(\bar{g},\psi)$. Using this, it is easy to see that the following is an isomorphism of topological groups
$P(\bar{f},\phi)\colon P(\overline{A})\rightleftarrows P(\overline{B}):P(\bar{g},\psi).$
For the converse direction, consider an isomorphism of procountable groups:
$\gamma:G\rightleftarrows H:\delta.$
Using the fact that the functor $P$ is full, we can rewrite this as follows (writing $\overline{A}$ in place of $(\overline{A},\overline{p})$ etc.)
$P(\bar{f},\phi)\colon P(\overline{A})\rightleftarrows P(\overline{B}):P(\bar{g},\psi),$
where $G=\colon P(\overline{A}),H=P(\overline{B})$ and $P(\bar{f},\phi)=\gamma,P(\bar{g},\psi)=\delta$. It is easy to verify that
$(\bar{f},\phi)\colon\overline{A}\rightleftarrows\overline{B}:(\bar{g},\psi)$
is an isomorphism in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. ∎
4.2. Proof of 4.1
Let the maps $p_{n}\colon T_{n+1}\to T_{n}$ and $q_{n}\colon U_{n+1}\to U_{n}$ be given by the parent relation on the trees $T$ and $U$, respectively. This yields inverse sequences $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{EpiGrp}$. Let $V_{n}=L(T_{n})$ and $W_{n}=L(U_{n})$.
First suppose that $T]\cong_{u}[U]$ via $\Phi$. By 3.5 there are increasing functions $\phi,\psi$ on $\mathbb{N}$ so that for each $x\in[T]$ and $y\in[U]$ we have
$\Phi(x)\upharpoonright n=r_{n}(x\upharpoonright{\phi(n)})$ and $\Phi^{-1}(y)\upharpoonright n=s_{n}(y\upharpoonright{\psi(n)})$,
for some onto functions $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$.
We induce morphisms in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$ by pre-morphisms
$(\bar{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\bar{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$
where $f_{n}=\widehat{r}_{n}$, $q_{n}=\widehat{s}_{n}$. For each $t\in T_{\phi(\psi(n))}$ and $u\in U_{\psi(\phi(n))}$ we have
$(s_{n}\circ r_{\psi(n)})(t)\subseteq t\text{ and }(r_{n}\circ s_{\phi(n)})(u)\subseteq u.$
So $(V_{n},\widehat{p}_{n})_{n}$ and $(W_{n},\widehat{q}_{n})_{n}$ are isomorphic in the category $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. To see this, recall that the binding maps $\widehat{p}_{n}$ and $\widehat{q}_{n}$ act as “predecessor maps”, since so do the maps $p_{n}$ and $q_{n}$. Hence $G_{T}$ and $G_{U}$ are isomorphic by 4.4.
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 4.4 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. By virtue of being pre-morphisms, the following diagrams are commutative for each $k>n$:
$V_{\phi(k)}$$W_{k}$$V_{\phi(n)}$$W_{n}$$\scriptstyle{f_{k}}$$\scriptstyle{\widehat{p}_{\phi(k),\phi(n)}}$$\scriptstyle{\widehat{q}_{k,n}}$$\scriptstyle{f_{n}}$$W_{\psi(k)}$$V_{k}$$W_{\psi(n)}$$V_{n}$$\scriptstyle{g_{k}}$$\scriptstyle{\widehat{q}_{\psi(k),\psi(n)}}$$\scriptstyle{\widehat{p}_{k,n}}$$\scriptstyle{g_{n}}$
Because the pre-morphisms induce inverses of each other, we have
($\star$) $\widehat{p}_{\phi(\psi(n)),n}=g_{n}\circ f_{\psi(n)}\qquad\widehat{q}_{\psi(\phi(n)),n}=f_{n}\circ g_{\phi(n)}.$
Let $r_{n}\colon\subseteq T_{\phi(n)}\to U_{n}$ be defined by
$r_{n}(t)=u$ if $f_{n}(t)$ is conjugate in $W_{n}$ to $u$.
(Note here that $f_{n}(t)^{2}=e$, so $f_{n}(t)$ is conjugate to a unique generator, or it equals $e$. The second case could prima facie happen, so $r_{n}$ could be a partial map, as indicated in the notation.) Similarly, let $s_{n}\colon\subseteq U_{\psi(n)}\to T_{n}$ be defined by
$s_{n}(u)=t$ if $g_{n}(u)$ is conjugate in $V_{n}$ to $t$.
We claim that $s_{n}\circ r_{\psi(n)}=p_{\phi(\psi(n)),n}$ for each $n$. To verify this, let $t\in T_{\phi(\psi(n))}$. By ($\star$4.2) we have that $g_{n}(f_{\psi(n)}(t))\neq e$. So
$f_{\psi(n)}(t)\sim_{W_{\psi(n)}}r_{\psi(n)}(t)\in U_{\psi(n)}$
where $\sim_{L}$ denotes conjugacy in a group $L$. So by ($\star$4.2) again
$\widehat{p}_{\phi(\psi(n)),n}(t)=g_{n}(f_{\psi(n)}(t))\sim_{V_{n}}g_{n}(r_{\psi(n)}(t))\sim_{V_{n}}s_{n}(r_{\psi(n)}(t)).$
Similarly, $r_{n}\circ s_{\phi(n)}=q_{\psi(\phi(n)),n}$. Since the compositions are onto and total, each map $r_{n}$ and $s_{n}$ is onto. Thus the maps $r_{n}$ and $s_{n}$ are also total (composing a partial after an onto map yields a partial map).
Corresponding to the diagrams in EpiGrp above, we have commutative diagrams of onto maps between sets
$T_{\phi(k)}$$U_{k}$$T_{\phi(n)}$$U_{n}$$\scriptstyle{r_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{r_{n}}$$U_{\psi(k)}$$T_{k}$$U_{\psi(n)}$$T_{n}$$\scriptstyle{s_{k}}$$\scriptstyle{q_{\psi(k),\psi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{s_{n}}$
Hence $T]\cong_{u}[U]$ by 3.5, as required.