### v2

# Procountable groups are not classifiable by countable structures

Abstract.
We prove that topological isomorphism on $2$-nilpotent procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ that expresses that two sequences of reals have bounded difference is Borel reducible to it. This marks substantial progress on an open problem of [16]: to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. A. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-1931. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (consolidator grant) “Groups, Crystals and Classifications”.
1. Introduction
It is often useful to have concrete presentations of abstract mathematical objects of a certain type. In this case one is interested in the complexity of whether two presentations present the same abstract object. For instance, finite graphs can be concretely presented by adjacency matrices (written as binary strings). The graph isomorphism problem asks whether two such matrices present the same graph. This problem is in NP, and its exact complexity has been the object of much recent study [1, 31]. The triviality problem for groups is to determine whether a presentation given by finitely many generators and relators presents the trivial group. This is undecidable [30], so the $\Sigma^{0}_{1}$ relation of isomorphism between finite presentations of groups is undecidable. In fact, every $\Sigma^{0}_{1}$ equivalence relation is computably reducible to it by [23] (also see [26, Cor. 2.6]).
For many types of mathematical objects, a countably infinite amount of information is required to present a member. In this case, one can usually view the presentations as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in [7], is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation [8] denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
A programme of Kechris et al. A Polish group is a topological group with a Polish topology, namely, one that is separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups [2]. Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ [2]. It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space (see [16]). So we can take the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ as the concrete presentations of non-archimedean Polish groups.
It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space; see for instance Kechris et al. [16]. In that article they initiated the following programme:
Given a Borel class $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$, determine the complexity of topological isomorphism between groups in $\mathcal{C}$.
There has been some progress on this program. We summarize the results known to us. One class considered in [16] is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [7, (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in [24]; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see [25, 27]; in particular in [27] it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in [4], where it is shown that the setwise equality of infinite sequences of reals is Borel reducible to it. Work in progress of Ruiwen Li, in conjunction with [4], would show that in fact GI is reducible to it.
We note that the programme of [16] asked to first determine whether a given class of interest $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. Borelness is unknown for the class of topologically finitely generated groups, and known to fail for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be properly co-analytic by Malicki [19]. Paolini and Shelah [29] proved that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete co-analytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Mekler [22] (as noted in [7, 2.3]) established this for the countable groups that are nilpotent of class $2$, and of exponent a fixed odd prime. His method of coding graphs into such groups will be instrumental for our main result, Theorem 1.2. Paolini and Shelah [28] showed it for countable torsion-free abelian groups. Most recently, Gao and Li [10] did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
The main result. Our work marks considerable progress on an open problem of Kechris et al. [16] pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of 2-nilpotent procountable groups of exponent a fixed odd prime $p$ already has a complexity much higher than GI.
Definition 1.1.
A topological group $G$ is called procountable if $G$ is topologically isomorphic to the inverse limit of an inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ of countable groups and onto homomorphisms $p_{n}\colon G_{n+1}\to G_{n}$. The topology of $G$ is the subspace topology given by the product $\prod_{n}G_{n}$, where each $G_{n}$ is equipped with the discrete topology.
It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., [20]. We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [12, Theorem 3.3] or [20, Lemma 2]). The procountable groups satisfying some law (such as being 2-nilpotent) then also form a Borel set: from a closed subgroup $G$ of $\mathrm{Sym}(\mathbb{N})$, using [14, Th. 12.13] one can in a Borel way determine a countable dense subgroup $D$, and it satisfies to require that $D$ satisfies the law.
Since each countable discrete group is procountable, by the aforementioned result of Mekler [22], GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality of sequences of real numbers. By a celebrated theorem of Kechris and Louveau [15], $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation to which $E_{1}$ is Borel reducible has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
$(x_{n})\ell_{\infty}(y_{n})\iff\exists M\in\mathbb{R}\ \forall n\in\mathbb{N}\ |x_{n}-y_{n}|<M.$
By a result of Rosendal (see [8, Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
Theorem 1.2.
Fix an odd prime $p$. The equivalence relation $\ell_{\infty}$ is Borel reducible to the topological isomorphism relation on 2-nilpotent procountable groups of exponent $p$. Consequently, the topological isomorphism relation on such procountable groups is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, it is not Borel reducible to GI.
It is known that being Borel reducible to GI is equivalent to being below the orbit equivalence relation of a Borel action of $\mathrm{Sym}(\mathbb{N})$, which is also equivalent to being classifiable by some sort of countable structures. It follows from our main theorem that the topological isomorphism of procountable groups is not classifiable by countable structures.
The proof makes use of the notion of uniform homeomorphism.
Definition 1.3.
A Polish metric space is a metric space $(X,d)$ where $d$ is a complete metric and the topology generated by $d$ is Polish. Let $X$ and $Y$ be Polish metric spaces. A bijection $\Phi$ between $X$ and $Y$ is called a uniform homeomorphism if both $\Phi$ and its inverse are uniformly continuous. $X$ and $Y$ are said to be uniformly homeomorphic if there is a uniform homeomorphism between $X$ and $Y$.
The two steps of the proof of the main result 1.2.
  1. 1.
    Section 2 shows that $\ell_{\infty}$ is Borel reducible to the relation of uniform homeomorphism between path spaces $[T]$, where $T$ is a pruned subtree of $\mathbb{N}^{<\mathbb{N}}$ and $[T]$ carries the standard ultrametric.
  2. 2.
    Section 4 reduces this uniform homeomorphism relation to the topological isomorphism relation on the Borel space of procountable nilpotent groups of class $2$ and exponent $p$, for $p$ an odd prime.
The second step only uses a very particular type of groups that could be called “solenoid groups”. All the groups $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ are isomorphic. Each group $G_{n}$ is generated by a set $V_{n}$, and the binding maps $p_{n}$ are determined by surjections $V_{n+1}\to V_{n}$.
Open questions.
We close the introduction with some open questions. The first two will be discussed again in see Section 2.4. An analytic equivalence relation on a Borel space is called universal analytic if each analytic equivalence relation is Borel reducible to it.
Question 1.4.
Is the equivalence relation of topological isomorphism on procountable groups universal analytic?
We note that our original construction for the second step used as $G_{n}$ the free Coxeter group generated by $V_{n}$, namely, the group freely generated by involutions $g_{v}$ for $v\in V_{n}$; see the arxiv version 2512.12256.v1. That approach does not yield nilpotent-2 groups. However, it is somewhat simpler to verify, because the only elements of order 2 in $G_{n}$ are the conjugates of the generators, so that the elements of $V_{n}$ up to conjugacy can be recovered from $G_{n}$. It is possible that this method could lead to an affirmative answer Question 1.4.
It is known that the relation of uniform homeomorphism on Polish metric spaces is universal analytic [5].
Question 1.5.
Is the equivalence relation of uniform homeomorphism on ultrametric Polish spaces universal analytic?
By the second step above, an affirmative answer to the second question would also answer the first in the affirmative.
Question 1.6.
Are the abelian procountable groups classifiable by countable structures?
Notice that it follows from [28] that GI is a lower bound for the complexity of this classification problem.
2. Borel reduction of $\ell_{\infty}$ to uniform homeomorphism
Definition 2.1.
By a tree on $\mathbb{N}$ we mean a nonempty set $T\subseteq\mathbb{N}^{<\mathbb{N}}$ that is closed under initial segments. The path space
$T]=\big\{x\in\mathbb{N}^{\mathbb{N}}\colon\forall n\in\mathbb{N}\ (x\!\!\upharpoonright\!n\in T)\big\}$
is a Polish ultrametric space with the metric given by declaring for $x\neq y$ that
[…]
The elements of $[T]$ will be called branches of $T$.
A tree $T$ is called pruned if for any $s\in T$ there is $t\in T$ such that $s\subsetneq t$. The space of all pruned trees on $\mathbb{N}$ can be seen to be a standard Borel (even Polish) subspace of $2^{(\mathbb{N}^{<\mathbb{N}})}$, which will be denoted $\mathcal{T}$.
Note that the paths ultrametric spaces are in a canonical bijection with the pruned trees, so we can see the elements of $\mathcal{T}$ as codes for such spaces.
Definition 2.2.
By $\cong_{u}$ we denote the equivalence relation of uniform homeomorphism on $\mathcal{T}$.
The following quick observation gives a lower bound on the complexity of $\cong_{u}$.
Proposition 2.3.
GI is Borel reducible to $\cong_{u}$.
Proof.
We use the main result of [3], where it was shown that GI is Borel bireducible with the homeomorphism relation between 0-dimensional compact metric spaces. Let $\Phi$ denote a Borel witness to this reduction. Then for any graphs $\Gamma,\Gamma^{\prime}$ with vertex set $\mathbb{N}$, $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $\Phi(\Gamma)$ and $\Phi(\Gamma^{\prime})$ are homeomorphic. Note that a Polish space is 0-dimensional if and only if it admits a compatible ultrametric. So there exists a compatible ultrametric $d_{\Gamma}$ for $\Phi(\Gamma)$. By the standard tools from descriptive set theory, we may find a Borel map giving the correspondence $\Phi(\Gamma)\mapsto d_{\Gamma}$.
Note that for any compact ultrametric space $(X,d_{X})$, the set of nonzero distances $\{d_{X}(x,y)\colon x,y\in X\}\setminus\{0\}$ is a decreasing sequence of positive real numbers converging to 0. Up to homeomorphism, we may assume that this sequence is $\{2^{-n}\colon n\in\mathbb{N}\}$. Now that $(\Phi(\Gamma),d_{\Gamma})$ is a compact ultrametric space with its set of nonzero distances a subset of $\{2^{-n}\colon n\in\mathbb{N}\}$, there is a unique tree $T_{\Gamma}\in\mathcal{T}$ such that $(\Phi(\Gamma),d_{\Gamma})$ is isometric to $T_{\Gamma}]$. Again the map $d_{\Gamma}\mapsto T_{\Gamma}$ is Borel.
Finally, note that, as compact spaces, $T_{\Gamma}]$ and $T_{\Gamma^{\prime}}]$ are homeomorphic if and only if they are uniformly homeomorphic. Thus we have that $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $T_{\Gamma}]$ and $T_{\Gamma^{\prime}}]$ are uniformly homeomorphic as compact ultrametric spaces. This proves the proposition. ∎
It follows that $\cong_{u}$ is a complete analytic set and hence non-Borel.
The rest of this section will establish the following result of independent interest, which is the first step towards proving our Main Result 1.2.
Theorem 2.4.
$\ell_{\infty}$ is Borel reducible to $\cong_{u}$, namely, uniform homeomorphism of the ultrametric spaces denoted by points in $\mathcal{T}$. Consequently, $\cong_{u}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. In particular, $\cong_{u}$ is not Borel reducible to GI.
2.1. Some facts on uniform homeomorphisms
The following is a direct consequence of the definition 2.1 of the metric on path spaces, and the definition of uniform continuity.
Lemma 2.5.
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$. Then $\Phi$ is uniformly continuous if and only if for all $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that for any $x,y\in[T]$, if $x\!\!\upharpoonright\!m=y\!\!\upharpoonright\!m$, then $\Phi(x)\!\!\upharpoonright\!n=\Phi(y)\!\!\upharpoonright\!n$.
In this context we define $\Omega_{\Phi}(n)$ to be the least $m$ such that the condition in the above lemma holds. Note that the lemma implies
$\big|\big\{t\in T\colon|t|=\Omega_{\Phi}(n)\big\}\big|\geqslant\big|\big\{s\in S\colon|s|=n\big\}\big|.$
Lemma 2.6.
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be a uniform homeomorphism (see 1.3). If both $[T],[S]$ are uncountable, then $\Omega_{\Phi}$ and $\Omega_{\Phi^{-1}}$ are unbounded monotone functions from $\mathbb{N}$ to $\mathbb{N}$.
Proof.
The monotonicity is obvious. If $\Omega_{\Phi}$ is bounded by $m$, then for each $n$ the string $\Phi(x)\!\!\upharpoonright\!n$ is determined by $x\!\!\upharpoonright\!m$. So $[S]$ is countable. ∎
2.2. Definition of the reduction
We will need to restrict the entries of the sequences in $\mathbb{R}^{\mathbb{N}}$ to natural numbers. This does not decrease the complexity of the equivalence relation:
Lemma 2.7.
$\ell_{\infty}\leqslant_{B}\ell_{\infty}\!\!\upharpoonright\!\mathbb{N}^{\mathbb{N}}$.
Proof.
Define $f\colon\mathbb{R}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$ by
$f(x)(2n)=\left\{\begin{array}[]{ll}\lfloor x(n)\rfloor,&\mbox{if $x(n)\geqslant 0$,}\\ 0,&\mbox{otherwise,}\end{array}\right$.
and
$f(x)(2n+1)=\left\{\begin{array}[]{ll}0,&\mbox{if $x(n)\geqslant 0$,}\\ -\lfloor x(n)\rfloor,&\mbox{otherwise.}\end{array}\right$.
Then for any $x,y\in\mathbb{R}^{\mathbb{N}}$ and $n\in\mathbb{N}$, we have $|f(x)(n)-f(y)(n)|\leqslant|x(n)-y(n)|$. Thus, if $x\ell_{\infty}y$, then $f(x)\ell_{\infty}f(y)$. Conversely, if |f(x)(n)f(y)(n)|<M|f(x)(n)-f(y)(n)|<M for all $n\in\mathbb{N}$, then for all $n\in\mathbb{N}$, either $x(n)y(n)>0$ and |x(n)y(n)|<M|x(n)-y(n)|<M hold, or else $x(n)y(n)\leqslant 0$ and $|x(n)|,|y(n)|<M$ hold; in both cases we have |x(n)y(n)|<2M|x(n)-y(n)|<2M. Thus $f$ is a reduction from $\ell_{\infty}$ to $\ell_{\infty}\!\!\upharpoonright\!\mathbb{N}^{\mathbb{N}}$. ∎
For the next step of the reduction, we consider an arbitrary $x\in\mathbb{N}^{\mathbb{N}}$ and define a pruned tree $\mathsf{T}_{x}$ on $\mathbb{N}$. Before defining $\mathsf{T}_{x}$, we first introduce some auxiliary trees
$\mathsf{S}_{*},\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots,\mathsf{S}_{\omega},\mathsf{R}_{0},\mathsf{R}_{1},\dots,\mathsf{R}_{k},\dots$
that will be used as building blocks. Let $\mathsf{S}_{*}\subseteq\mathbb{N}^{<\mathbb{N}}$ be the tree consisting of all finite sequences $s\in\mathbb{N}^{<\mathbb{N}}$ such that $s(n)$ is even for each $n\in\mathbb{N}$. Then $\mathsf{S}_{*}$ is a copy of the full tree $\mathbb{N}^{<\mathbb{N}}$. The point of defining $\mathsf{S}_{*}$ is that we may arbitrarily extend $\mathsf{S}_{*}$ by adding extensions of existing nodes, and still maintain that the extended tree be an “official” subtree of $\mathbb{N}^{<\mathbb{N}}$.
The trees $\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots$ will be abstractly and inductively defined. Their exact realizations as subtrees of $\mathbb{N}^{<\mathbb{N}}$ will not be important. To begin, let $\mathsf{S}_{1}$ be a single branch. Thus the Cantor–Bendixson rank of $\mathsf{S}_{1}]$ is $1$. Let $\mathsf{S}_{2}$ be a pruned tree so that the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $\mathsf{S}_{1}]$. So the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $2$. In general, let $\mathsf{S}_{k+1}$ be a pruned tree whose Cantor–Bendixson derivative is $\mathsf{S}_{k}]$. Then the Cantor–Bendixson rank of $\mathsf{S}_{k+1}$ is $k+1$. This finishes the definitions of $\mathsf{S}_{k}$ for $1\leqslant k<\mathbb{N}$.
Next we define $\mathsf{S}_{\omega}$ to be the extension of $\mathsf{S}_{*}$ defined as follows. For each $k\in\mathbb{N}$, let $\mathsf{z}_{k}$ be the branch of $\mathsf{S}_{*}$ defined as
$\mathsf{z}_{k}(m)=\left\{\begin{array}[]{ll}2k,&\mbox{if $m=0$,}\\ 0,&\mbox{otherwise.}\end{array}\right$.
For any $k,m\in\mathbb{N}$, let $\mathsf{t}_{k,m}=\mathsf{z}_{k}\!\!\upharpoonright\!(m+1)$. $\mathsf{S}_{\omega}$ is obtained by extending $\mathsf{S}_{*}$ with appending a copy of $\mathsf{S}_{k}$ at the node $\mathsf{t}_{k,2m}$ for all $k,m\in\mathbb{N}$.
The distinctive feature of $\mathsf{S}_{\omega}$ is that any autohomeomorphism of $\mathsf{S}_{\omega}]$ must keep $\mathsf{z}_{k}$ fixed for all $k\in\mathbb{N}$. This is because of the topological characterization of $\mathsf{z}_{k}$ as the only $z\in[\mathsf{S}_{\omega}]$ with the property that:
for any clopen set $U$ with $z\in U$, the $k$-th Cantor–Bendixson derivative of $U$ is the perfect kernel of $U$, and for any $i<k$, the $i$-th Cantor–Bendixson derivative of $U$ contains a perfect subset but is not itself perfect.
Next we abstractly define a sequence of finitely splitting pruned trees
$\mathsf{R}_{0},\mathsf{R}_{1},\dots,\mathsf{R}_{k},\dots$
so that no $\mathsf{R}_{k}]$ has an isolated point, as follows. For each $k\in\mathbb{N}$, let $\mathsf{R}_{k}$ be $2^{k}$-splitting on the first level, and then each node on the first level be extended with a copy of the full binary tree. It is easy to see that for any $k\in\mathbb{N}$, $\mathsf{R}_{k}]$ is homeomorphic to a Cantor set, and therefore there is a uniform homeomorphism between $\mathsf{R}_{k}]$ and $\mathsf{R}_{\ell}]$ for any $k,\ell\in\mathbb{N}$.
Finally we are ready to define $\mathsf{T}_{x}$ given an $x\in\mathbb{N}^{\mathbb{N}}$. $\mathsf{T}_{x}$ is obtained from $\mathsf{S}_{\omega}$ by extending the node $\mathsf{t}_{k,2m+1}$ by a copy of $\mathsf{R}_{x(k)}$ for all $k,m\in\mathbb{N}$. This finishes the definition of $\mathsf{T}_{x}$.
For each $x\in\mathbb{N}^{\mathbb{N}}$, we let $\mathsf{S}_{*}(\mathsf{T}_{x})$, $\mathsf{S}_{\omega}(\mathsf{T}_{x})$, $\mathsf{t}_{k,m}(\mathsf{T}_{x})$, $\mathsf{z}_{k}(\mathsf{T}_{x})$ etc. denote respectively the subtrees $\mathsf{S}_{*}$, $\mathsf{S}_{\omega}$, nodes $\mathsf{t}_{k,m}$, and branches $\mathsf{z}_{k}$ etc. in the construction of $\mathsf{T}_{x}$. Note that for different $x$ these objects are isomorphic, but we need to address the specific copies of them as substructures of $\mathsf{T}_{x}$.
For $k,m\in\mathbb{N}$, we let $\mathsf{R}_{x(k),m}$ denote the copy of $\mathsf{R}_{x(k)}$ which was appended at the node $\mathsf{t}_{k,2m+1}(\mathsf{T}_{x})$.
Note that for any $x\in\mathbb{N}^{\mathbb{N}}$ and $k\in\mathbb{N}$, $\mathsf{z}_{k}(\mathsf{T}_{x})$ is still topologically characterized by the same property as above, since the additional sets added beyond $\mathsf{S}_{\omega}]$ in the construction of $\mathsf{T}_{x}]$ are all perfect.
Therefore, if $\theta$ is any homeomorphism from $\mathsf{T}_{x}]$ to $\mathsf{T}_{y}]$, then we must have that $\theta(\mathsf{z}_{k}(\mathsf{T}_{x}))=\mathsf{z}_{k}(\mathsf{T}_{y})$ for all $k\in\mathbb{N}$.
Although we do not need this in the proof, we observe that for any $x,y\in\mathbb{N}^{\mathbb{N}}$, $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$ are homeomorphic.
Definition 2.8.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be Polish metric spaces and let $f\colon X\to Y$. Let $L>0$. We say that $f$ is $L$-Lipschitz if for any $x,x^{\prime}\in X$,
$d_{Y}(f(x),f(x^{\prime}))\leqslant L\,d_{X}(x,x^{\prime}).$
$f$ is Lipschitz if it is $L$-Lipschitz for some $L>0$.
For $k,\ell\in\mathbb{N}$, consider the natural homeomorphism $\Phi_{k,\ell}$ between $\mathsf{R}_{k}]$ and $\mathsf{R}_{\ell}]$. For later use we note that $\Phi_{k,\ell}$ is $L$-Lipschitz with $L=2^{\ell-k}$. Similarly, $\Phi_{k,\ell}^{-1}=\Phi_{\ell,k}$ is $L^{-1}$-Lipschitz.
2.3. Why the reduction works
Assume $x,y\in\mathbb{N}^{\mathbb{N}}$ with $x\ell_{\infty}y$. Let $M>0$ be such that |x(n)y(n)|<M|x(n)-y(n)|<M for all $n\in\mathbb{N}$. We define a uniform homeomorphism $\Phi$ between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$. For this, note that $\mathsf{T}_{x}]$ is the disjoint union of $\mathsf{S}_{\omega}(\mathsf{T}_{x})]$ and $\mathsf{R}_{x(k),m}]$ for all $m\in\mathbb{N}$, and $\mathsf{T}_{y}]$ is the disjoint union of $\mathsf{S}_{\omega}(\mathsf{T}_{y})]$ and $\mathsf{R}_{y(k),m}]$ for all $m\in\mathbb{N}$. Let $\Phi$ be the homeomorphism from $\mathsf{S}_{\omega}(\mathsf{T}_{x})]$ onto $\mathsf{S}_{\omega}(\mathsf{T}_{y})]$ induced by the identity map from $\mathsf{S}_{\omega}(\mathsf{T}_{x})$ to $\mathsf{S}_{\omega}(\mathsf{T}_{y})$. Then it is obvious that $\Phi$ is an isometry, and therefore both $\Phi$ and $\Phi^{-1}$ are $1$-Lipschitz. Let $\psi_{k,m}$ be the natural homeomorphism between $\mathsf{R}_{x(k),m}]$ and $\mathsf{R}_{y(k),m}]$. Then for all $k,m\in\mathbb{N}$, both $\psi_{k,m}$ and $\psi_{k,m}^{-1}$ are $2^{M}$-Lipschitz. Let $\Psi=\Phi\cup\bigcup_{k,m}\psi_{k,m}$. Then both $\Psi$ and $\Psi^{-1}$ are $2^{M}$-Lipschitz. Thus $\Psi$ is a uniform homeomorphism between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$.
Conversely, assume $\Psi$ is a uniform homeomorphism between $\mathsf{T}_{x}]$ and $\mathsf{T}_{y}]$. In particular, $\Psi$ is a homeomorphism. Then $\Psi(\mathsf{z}_{k}(\mathsf{T}_{x}))=\mathsf{z}_{k}(\mathsf{T}_{y})$ for any $k\in\mathbb{N}$. By Lemma 2.6, there is $m>1$ such that
$\Omega_{\Psi}(2m+2)>1$
and, letting $n=\Omega_{\Psi}(2m+2)$, also
$\Omega_{\Psi^{-1}}(2n+2)>1.$
Let $M$ be the least such that $2M+2\geqslant\Omega_{\Psi^{-1}}(2n+2)$. We claim that for any $k\in\mathbb{N}$, $\Psi([\mathsf{R}_{x(k),n}])$ is contained in
$\bigcup_{m<p\leqslant M}[\mathsf{R}_{y(k),p}].$
To see this, let $u\in[\mathsf{R}_{x(k),n}]$ be arbitrary. Then
$u\upharpoonright(2n+1)=\mathsf{z}_{k}(\mathsf{T}_{x})\upharpoonright(2n+1).$
Since $2n+1>n=\Omega_{\Psi}(2m+2)$, we have that
$\Psi(u)\upharpoonright(2m+2)=\mathsf{z}_{k}(\mathsf{T}_{y})\upharpoonright(2m+2).$
However, if $\Psi(u)=v\in[\mathsf{R}_{y(k),p}]$ for $p>M$, then
$v\upharpoonright(2p+1)=\mathsf{z}_{k}(\mathsf{T}_{y})\upharpoonright(2p+1),$
and since $2p+1>2M+2\geqslant\Omega_{\Psi^{-1}}(2n+2)$, we have
$u\upharpoonright(2n+2)=\mathsf{z}_{k}(\mathsf{T}_{x})\upharpoonright(2n+2),$
which is a contradiction.
By our construction, the number of nodes at level $2n+2$ in the copy of $\mathsf{R}_{x(k),n}$ is $2^{x(k)}$. The number of nodes at level $2M+2$ in the copies of $\bigcup_{m<p\leqslant M}\mathsf{R}_{y(k),p}$ is no more than
$2^{y(k)+(M-m)}+2^{y(k)+(M-m-1)}+\cdots+2^{y(k)}=2^{y(k)}(2^{M-m+1}-1)\leqslant 2^{y(k)+M-m+1}.$
By Lemma 2.5 we have that $2^{y(k)+M-m+1}\geqslant 2^{x(k)}$, and hence $x(k)-y(k)\leqslant M-m+1$. This means that x(k)y(k)x(k)-y(k) is bounded. By symmetry, y(k)x(k)y(k)-x(k) is also bounded. Therefore $x\ell_{\infty}y$.
2.4. Further remarks
Definition 2.9.
We call an analytic equivalence relation universal if any analytic equivalence relation is Borel reducible to it.
Such equivalence relations were called complete analytic equivalence relations in [5]. Here we call them universal to avoid possible confusion. In [5] and subsequent research, many natural equivalence relations were found to be universal analytic equivalence relations. In particular, it was shown in [5] that the uniform homeomorphism for all Polish metric spaces is a universal analytic equivalence relation.
We remark that our result does not show that $\cong_{u}$ is universal. By 2.3 GI is reducible to $\cong_{u}$, and by the main result 1.2 $\ell_{\infty}$ is reducible to $\cong_{u}$. However, the equivalence relation $c_{0}$ does not reduce to $\mbox{\rm GI}\times\ell_{\infty}$. Recall that $c_{0}$ is the equivalence relation on $\mathbb{R}^{\mathbb{N}}$ defined as
$(x_{n})c_{0}(y_{n})\iff\lim_{n}|x_{n}-y_{n}|=0.$
Toward a contradiction, assume that there is a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [13, Theorem 3.18 and Example 3.23]; also see [8, Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [8, Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [8, Theorem 8.5.2] and its proof).
We conjecture that $\cong_{u}$, and hence uniform homeomorphism of Polish ultrametric spaces, are universal analytic equivalence relations.
3. Preliminaries on procountable groups
In 1.1 we defined procountable groups. This section provides material on procountable groups that will be needed for the second step in our overall proof of the Main Result 1.2.
3.1. Presentations of procountable groups
Encoding inverse systems by structures with domain $\mathbb{N}$ that have infinitely many sorts, we obtain a standard Borel space $\mathcal{P}$ of presentations of procountable groups. We show that these presentations can be converted in a Borel way into procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$, and that the latter form a Borel set $\mathcal{C}$.
Notation 3.1.
The space of all finite sequences of natural numbers is denoted as $\mathbb{N}^{<\mathbb{N}}$. For $s,t\in\mathbb{N}^{<\mathbb{N}}$ and $x\in\mathbb{N}^{\mathbb{N}}$, we use $s\subseteq t$ to denote that $s$ is an initial segment of $t$ and similarly for $s\subseteq x$. For $s\in\mathbb{N}^{<\mathbb{N}}$, we write
$s]=\left\{x\in\mathrm{Sym}(\mathbb{N})\colon\,s\subseteq x\right\}.$
For a closed subgroup $G\leqslant_{c}\mathrm{Sym}(\mathbb{N})$, we will write $s]_{G}$ for $s]\cap G$. We let $\sigma_{n}$ denote the tuple $(0,1,\ldots,n-1)$; note that the $\sigma_{n}]_{G}$ form a neighbourhood basis of the identity in $G$ consisting of open subgroups.
Given $s,t\in\mathbb{N}^{<\mathbb{N}}$, let $s\circ t$ be the sequence representing composition, on the longest initial segment of $\mathbb{N}$ in which it is defined. Let $s^{-1}$ be the inverse of $s$ on the longest initial segment in which it is defined. For instance, $(7,4,3,1,0)\circ(3,4,6)=(1,0)$, and $(1,2,0,5)^{-1}=(2,0,1)$.
[…]
$G=\left\{x\in\prod_{n}G_{n}\colon\,\forall n\,[p_{n}(x(n+1))=x(n)]\right\}.$
Remark 3.2.
From an inverse system as above, one can in a Borel way determine a corresponding closed subgroup of $\mathrm{Sym}(\mathbb{N})$. One replaces $G_{n}$ by an isomorphic group with domain $D_{n}=\{\langle i,n\rangle\colon i\in\mathbb{N}\}$, where $\langle\cdot,\cdot\rangle\colon\mathbb{N}\times\mathbb{N}\to$$\mathbb{N}$ is the Cantor pairing function. Then the left regular presentation of $G_{n}$ is a copy of $G_{n}$ as a permutation group with support $D_{n}$, and the product of these copies in $\mathrm{Sym}(\mathbb{N})$ is a copy $C$ of $\prod_{n}G_{n}$ as a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. Finally, one obtains a copy of $G$ as a closed subgroup of $C$ in a Borel way from the binding maps $p_{n}$.
Proposition 3.3.
The procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel class $\mathcal{C}$.
Proof.
As mentioned before, $G$ is procountable if and only if the identity has a neighbourhood basis of normal open subgroups. In the setting of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ this is equivalent to: for each $n$, there is $k>n$ such that the normal closure (in $G$) of $\sigma_{k}]_{G}$ is contained in $\sigma_{n}]_{G}$. Since $\sigma_{n}]_{G}$ is a subgroup, it suffices to express as a Borel condition about $G$ that
(1) $\forall x\in G\,\left[x^{-1}\circ([\sigma_{k}]_{G})\circ x\subseteq[\sigma_{n}]_{G}\right].$
In the following, letters $s,t$ etc. range over sequences that have an extension in $G$. Recalling 3.1, we claim that (1) is equivalent to the Borel condition
(2) $\forall s\ \forall t\supseteq\sigma_{k}\ \ [s^{-1}\circ t\circ s]_{G}\cap[\sigma_{n}]_{G}\neq\emptyset.$
For (1) $\to$ (2), given $s$ and $t\supseteq\sigma_{k}$, let $x\supseteq s$, $x\in G$ and $y\supseteq t$, $y\in G$. Then $x^{-1}\circ y\circ x\in[\sigma_{n}]_{G}$.
For (2) $\to$ (1), suppose that there is $x\in G$ such that $x^{-1}\circ([\sigma_{k}]_{G})\circ x$ is not contained in $\sigma_{n}]_{G}$. Then, by the continuity of the conjugation operation, for sufficiently long $s\subseteq x$ and sufficiently long $t\supseteq\sigma_{k}$, we have $s^{-1}\circ t\circ s]_{G}\cap[\sigma_{n}]_{G}=\emptyset$. ∎
Remark 3.4.
[…]
3.2. Proω categories
We provide some preliminaries on the correspondence of isomorphisms of procountable groups and certain isomorphisms of inverse systems of type $\omega$. Our purpose is mainly to lay down a notational framework; more general results are obtained in [21] using a somewhat different language.
Definition 3.5.
Let $\mathcal{K}$ be a category that has as objects countable structures in the same signature, and as morphisms surjective homomorphisms between them. The category $\mathrm{Pro}_{\omega}(\mathcal{K})$ has as objects the inverse systems $(\overline{A},\overline{p})=(A_{n},p_{n})_{n\in\mathbb{N}}$ over $\mathcal{K}$, where $p_{n}\colon A_{n+1}\to A_{n}$. The pre-morphisms
$\langle\overline{f},\phi\rangle\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$
are given as sequences of morphisms $f_{n}:A_{\phi(n)}\to B_{n}$ where $\phi:\mathbb{N}\rightarrow\mathbb{N}$ is increasing, and all the diagrams commute. In more detail, for $k>n$, write $p_{k,n}$ for the “interval composition” maps $p_{n}\circ p_{n+1}\circ\ldots\circ p_{k-1}\colon A_{k}\to A_{n}$ and define $q_{k,n}$ in a similar way; we require that the diagrams
$A_{\phi(k)}$$B_{k}$$A_{\phi(n)}$$B_{n}$$\scriptstyle{f_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{f_{n}}$
commute. Composition of pre-morphisms is defined in the canonical way (see [21, pg. 5]), and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=\{{\mathrm{id}_{A_{n}}}\}_{n\in\mathbb{N}}$.
For each pair of objects $(\overline{A},\overline{p})$ and $(\overline{B},\overline{q})$ one defines an equivalence relation $\sim$ between pre-morphisms $(\bar{f},\phi),(\bar{f}^{\prime},\phi^{\prime}):(\overline{A},\overline{p})\to(\overline{B},\overline{q})$: let $(\bar{f},\phi)\sim(\bar{f}^{\prime},\phi^{\prime})$ if for each $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that $m\geqslant\phi(n),\phi^{\prime}(n)$ and the following diagram commutes:
$A_{m}$$A_{\phi(n)}$$A_{\phi^{\prime}(n)}$$B_{n}$$\scriptstyle{p_{m,\phi(n)}}$$\scriptstyle{p_{m,\phi^{\prime}(n)}}$$\scriptstyle{f_{n}}$$\scriptstyle{f^{\prime}_{n}}$
The transitivity of the equivalence relation $\sim$ and its compatibility with composition of pre-morphisms are proved in [21, pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are defined as the equivalence classes of pre-morphisms under $\sim$.
Remark 3.6.
Notice that for a pre-morphism $(\bar{f},\phi):(\overline{A},\overline{p})\to(\overline{A},\overline{p})$, we have $(\bar{f},\phi)\sim({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$ if and only if for every $n\in\mathbb{N}$ there is $m\geqslant\phi(n),n$ such that the diagram above commutes where $\phi^{\prime}(n)=n$, $B_{n}=A_{n}$ and $f_{n}^{\prime}$ is the identity of $A_{n}$. Since the morphisms of $\mathcal{K}$ are surjections, this simply means that $f_{n}=p_{\phi(n),n}$ for each $n$.
Let $\mathrm{GrpEpi}$ be the category of countable groups with onto homomorphisms.
Lemma 3.7.
  1. (i)
    The inverse limit operation yields a full functor $P$ from $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$ to the category of procountable groups with continuous homomorphisms that have dense range.
  2. (ii)
    $P$ preserves isomorphisms in both directions.
Proof.
All inverse systems will be of type $\omega$.
[…]
In order to show that the functor $P$ is full, suppose that $\gamma\colon G\to H$ is a continuous homomorphism with dense range. We will define a pre-morphism $(\bar{f},\phi)\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$ with $P(\bar{f},\phi)=\gamma$. Let $a_{n}$ and $b_{n}$ be the kernels of the canonical projections $\pi^{G}_{n}:G\to A_{n}$ and $\pi^{H}_{n}:H\to B_{n}$ respectively. Let $\phi$ be an increasing function such that $a_{\phi(n)}\subseteq\gamma^{-1}(b_{n})$. Such $\phi$ exists because $\gamma^{-1}({b_{n}})$ is an open subgroup of $A_{n}$, and the $a_{n}$} form a basis of neighbourhoods of $e$. (This part of the argument is special to procountable groups.) For $n\in\mathbb{N}$ we define $f_{n}:A_{\phi(n)}\rightarrow B_{n}$ to be the unique map such that the diagram
[…]
commutes. Since $\gamma$ has dense range, $\pi^{H}_{n}\circ\gamma$ is onto, and hence $f_{n}$ is onto. To verify that for each $k>n$ the diagram
$A_{\phi(k)}$$B_{k}$$A_{\phi(n)}$$B_{n}\ .$$\scriptstyle{f_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{f_{n}}$
commutes, it suffices to consider the diagram
[…]
and realise that the upper rectangle and the outer rectangle both commute, noting that $p_{\phi(k),\phi(n)}\circ\pi^{G}_{\phi(k)}=\pi^{G}_{\phi(n)}$ and $q_{k,n}\circ\pi^{H}_{k}=\pi^{H}_{n}$.
(ii). For the forward direction, consider an isomorphism
$(\bar{f},\phi)\colon\overline{A}\rightleftarrows\overline{B}:(\bar{g},\psi)$
in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall to equal
$A_{\phi(\psi(n))}\xrightarrow{f_{\psi(n)}}B_{\psi(n)}\xrightarrow{g_{n}}A_{n},$
is such that $(\bar{g},\psi)\circ(\bar{f},\phi)\sim(\mathrm{id}_{\overline{A}},\mathrm{id}_{\mathbb{N}})$. By 3.6, this simply means that $g_{n}\circ f_{\psi(n)}=p_{\phi(\psi(n)),n}$. We argue similarly for the composition $(\bar{f},\phi)\circ(\bar{g},\psi)$. Using this, it is easy to see that the following is an isomorphism of topological groups
$P(\bar{f},\phi)\colon P(\overline{A})\rightleftarrows P(\overline{B}):P(\bar{g},\psi).$
For the converse direction, consider an isomorphism of procountable groups:
$\gamma:G\rightleftarrows H:\delta.$
Using the fact that the functor $P$ is full, we can rewrite this as follows (writing $\overline{A}$ in place of $(\overline{A},\overline{p})$ etc.)
$P(\bar{f},\phi)\colon P(\overline{A})\rightleftarrows P(\overline{B}):P(\bar{g},\psi),$
where $G=\colon P(\overline{A}),H=P(\overline{B})$ and $P(\bar{f},\phi)=\gamma,P(\bar{g},\psi)=\delta$. It is easy to verify that
$(\bar{f},\phi)\colon\overline{A}\rightleftarrows\overline{B}:(\bar{g},\psi)$
is an isomorphism in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. ∎
4. Borel reduction of uniform homeomorphism to isomorphism of 2-nilpotent procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, as laid out in the introduction. In 4.7 we will exhibit a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of nilpotent-2 procountable groups of exponent $p$, where $p$ is an odd prime.
4.1. The 2-nilpotent groups $L(X)$
Let $N$ be the free nilpotent group of class $2$ and exponent $p$ on free generators $\{x_{i}:i\in\mathbb{N}\}$. For distinct $r,s\in\mathbb{N}$ we write
$x_{r,s}=[x_{r},x_{s}].$
As noted in Mekler [22], the centre $Z(N)$ of $N$ is an elementary abelian $p$-group (and hence an $\mathbb{F}_{p}$ vector space) with basis $x_{r,s}$ for $r<s$. Given an undirected irreflexive graph $A$ with vertex set $\mathbb{N}$, Mekler sets
($\diamond$) $G(A)=N/\langle x_{r,s}\colon\,rAs\rangle.$
The centre $Z=Z(G(A))$ is an elementary abelian $p$-group with basis the $x_{r,s}$ such that $\lnot rAs$. Also $G(A)/Z$ is an abelian group of exponent $p$ freely generated by the $Zx_{i}$ (recall that in a $2$-nilpotent group the commutator subgroup is contained in the center). Mekler obtained a normal form for elements of $G(A)$.
Lemma 4.1.
(i) Every element $c$ of $Z$ can be uniquely written in the form
$\prod_{(r,s)\in E}x_{r,s}^{\beta_{r,s}}$
where $E\subseteq\omega\times\mathbb{N}$ is a finite set of pairs $(r,s)$ with $r<s$ and $\lnot rAs$, and $0<\beta_{r,s}<p$.
(ii) Every element of $G(A)$ can be uniquely written in the form $c\cdot v$ where $c\in Z$, and $v=\prod_{i\in C}x_{i}^{\alpha_{i}}$, for $C\subseteq\mathbb{N}$ finite and $0<\alpha_{i}<p$. (The product $\prod_{i\in C}x_{i}^{\alpha_{i}}$ is interpreted along the indices in ascending order.)
The following fact already noted in [22] is evident.
Fact 4.2.
Let $C,D\subseteq\mathbb{N}$. The following holds in $G(A)$.
$\prod_{r\in C}{x_{r}}^{{\alpha}_{r}},\prod_{s\in D}{x_{s}}^{{\beta}_{s}}]=\prod_{r\in C,\,s\in D,\,r<s,\,\lnot Ars}x_{r,s}^{\alpha_{r}\beta_{s}-\alpha_{s}\beta_{r}}$
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2. (While we will no use it, we note that this graph is “nice” in the sense of Mekler, cf. [22, Definition 1.1].)
Notation 4.3.
For $n\in\mathbb{N}$ write $n^{*}=\lfloor n/2\rfloor$. Let $A=\{\langle r,s\rangle\colon\,r\neq s\land r^{*}=s^{*}\}$. Let $L=G(A)$. For $u\in L$ we write $\underline{u}=uZ(L)$, which is an element of the group $L/Z(L)$.
The idea is that $G(A)$ has the minimum commutativity: any pair of elements that commute nontrivially have to generate the same subgroup as some pair $x_{v},x_{v+1}$ for even $v$, all modulo the centre.
Lemma 4.4.
Suppose that $c,d\in L$ are such that $[c,d]=e$.
Either there is an even $i\in\mathbb{N}$ such that $\langle\underline{c},\underline{d}\rangle=\langle\underline{x}_{i},{\underline{x}}_{i+1}\rangle$, or the rank of $\langle\underline{c},\underline{d}\rangle$ as a subgroup of $L/Z(L)$ is at most $1$.
Proof.
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 4.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose the first condition fails. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption’ there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 4.2.
Next we claim that $C=D$. Assume for contradiction that $C$ is not contained in $D$ (the case that $D$ is not contained in $C$ is symmetric). Let $r\in C\setminus D$. Then there is $s\in D$ such that $r^{*}\neq s^{*}$ because $D$ is not a singleton. So $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$. Thus $C=D$.
Now let $m=\min(C)$, and let $\lambda=\alpha_{m}/\beta_{m}$ as an element of $\mathbb{F}_{p}$. If $k\in C$ and $m^{*}<k^{*}$, then $\alpha_{k}/\beta_{k}=\lambda$. Some such $k$ exists. So if $m$ is even and $m+1\in C$, then we also have $\alpha_{m+1}/\beta_{m+1}=\lambda$.
We conclude that $\alpha_{r}=\beta_{r}\lambda$ holds in $\mathbb{F}_{p}$ for each $r\in C$. This shows that $\underline{c}=\underline{d}^{\lambda}$, and hence that the rank of $\langle\underline{c},\underline{d}\rangle$ is at most $1$. ∎
Definition 4.5.
Given a countably infinite set $X$, let $N(X)$ be the free 2-nilpotent group of exponent $p$ on distinct free generators $a_{v},b_{v}$ ($v\in X$). Let
$L(X)=N(X)/\langle\{[a_{v},b_{v}]\colon v\in X\}\rangle.$
Given a set $X$, for $v\in X$ we will interpret $a_{v}$ and $b_{v}$ in $L(X)$. Any surjection $q\colon X\to Y$ induces a group epimorphism $\widehat{q}\colon L(X)\to L(Y)$ given by mapping the equivalence class of $a_{v}$ to the equivalence class $a_{q(v)}$, and similarly for $b_{v}$ and $b_{q(v)}$.
An arbitrary bijection $h\colon X\to\mathbb{N}$ induces an isomorphism $L(X)\to L$ (for $L$ as in 4.3) by sending $a_{v}$ to the equivalence class of $x_{2h(v)}$ and $b_{v}$ to the equivalence class of $x_{2h(v)+1}$ given by the presentation in ($\diamond$4.1). So we can apply Lemma 4.4 to $L(X)$ as well. This lemma can now be used to turn certain group epimorphism $L(X)\to L(Y)$ into surjections $X\to Y$ in a functorial manner. This will be needed to eventually verify that the reduction for the second step (4.7 below) works.
If $r\colon G\to H$ is an epimorphism of groups, we denote by $\underline{r}$ the induced epimorphism $G/Z(G)\to H/Z(H)$.
Lemma 4.6.
Let $\mathcal{L}$ be the category of groups of the form $L(X)$ with morphisms the epimorphism $g\colon L(X)\to L(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $L(Y)/Z(L(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with surjections is declared by sending $L(X)$ to $X$, and sending a morphism $g:L(X)\to L(Y)$ to the map $p=F(g)\colon X\to Y$ such that
$\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle=\langle\underline{a_{p(v)}},\underline{b_{p(v)}}\rangle.$
Furthermore, if $q\colon X\to Y$ is a surjection, then $\widehat{q}$ is a morphism of $\mathcal{L}$, and $F(\widehat{q})=q$.
Proof.
To show that $F$ is well-defined for morphisms of $\mathcal{L}$, it suffices to apply 4.4 to $L=L(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. One verifies easily that $F$ is compatible with composition, and that $F(\widehat{q})=q$. ∎
4.2. The reduction for the second step
Definition 4.7.
Given a pruned tree $T$, let $T_{n}=\{s\in T\colon|s|=n\}$ be the $n$-th level of the tree $T$, and let $p_{n}\colon T_{n+1}\to T_{n}$ be the predecessor map. We may assume that $T_{1}$ is infinite because the trees we obtain through the coding in 2.4 have that property. Let
[…]
We view $G_{T}$ concretely as a closed subgroup of $\mathrm{Sym}(\mathbb{N})$ according to Remark 3.2. Needless to say, the map sending $T$ to $G_{T}$ is Borel. The rest of this section will establish the following, which is the second step towards proving the Main Result 1.2.
Theorem 4.8.
Let $T$ and $U$ be pruned trees on $\mathbb{N}$ such that $T_{1}$ and $U_{1}$ are infinite. Then $T]\cong_{u}[U]$ if and only if $G_{T}$ is topologically isomorphic to $G_{U}$.
Proof.
Let the maps $p_{n}\colon T_{n+1}\to T_{n}$ and $q_{n}\colon U_{n+1}\to U_{n}$ be given by the predecessor maps on the trees $T$ and $U$, respectively. Write
$V_{n}=L(T_{n})\qquad W_{n}=L(U_{n}).$
So we have inverse systems $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{GrpEpi}$ determining $G_{T}$ and $G_{U}$, respectively (see 4.5).
First suppose that $T]\cong_{u}[U]$ via a uniform homeomorphism $\Phi$. By 2.5 there are increasing functions $\phi,\psi$ on $\mathbb{N}$ and onto functions
$r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$
so that for each $x\in[T]$ and $y\in[U]$ we have
$\Phi(x)\upharpoonright n=r_{n}(x\upharpoonright{\phi(n)})$ and $\Phi^{-1}(y)\upharpoonright n=s_{n}(y\upharpoonright{\psi(n)})$.
Let $f_{n}=\widehat{r}_{n}$ and $q_{n}=\widehat{s}_{n}$ as defined in 4.5. The pre-morphisms
$(\bar{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\bar{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$
induce morphisms in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. For each $t\in T_{\phi(\psi(n))}$ and $u\in U_{\psi(\phi(n))}$,
$(s_{n}\circ r_{\psi(n)})(t)\subseteq t\text{ and }(r_{n}\circ s_{\phi(n)})(u)\subseteq u.$
So $(V_{n},\widehat{p}_{n})_{n}$ and $(W_{n},\widehat{q}_{n})_{n}$ are isomorphic in the category $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$ (to see this, recall that the binding maps $\widehat{p}_{n}$ and $\widehat{q}_{n}$ act as “predecessor maps”, since so do the maps $p_{n}$ and $q_{n}$). Hence $G_{T}$ and $G_{U}$ are isomorphic by 3.7.
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 3.7 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Because the pre-morphisms induce inverses of each other, we have
($\star$) $\widehat{p}_{\phi(\psi(n)),n}=g_{n}\circ f_{\psi(n)}\qquad\widehat{q}_{\psi(\phi(n)),n}=f_{n}\circ g_{\phi(n)}.$
Using the notation of 4.6, the maps on the left hand sides of the equation are morphisms of the category $\mathcal{L}$. We claim that $g_{n}$ and $f_{n}$ are also morphisms of $\mathcal{L}$ for each $n$. We verify it for $g_{n}$, the case of $f_{n}$ being symmetric. We use the left equation in ($\star$4.2). Write $p$ for $p_{\phi(\psi(n)),n}$. For each $v\in T_{\phi(\psi(n)}$, by definition we have $\widehat{p}(\langle a_{v},b_{v}\rangle)=\langle a_{p(v)},b_{p(v)}\rangle$. Thus the rank of $\langle\underline{a}_{v},\underline{b}_{v}\rangle$ cannot be reduced by applying the homomorphism induced by $f_{\psi(n)}$ on the central quotients. So $f_{\psi(n)}$ is in $\mathcal{L}$ by 4.6, and hence $g_{n}$ is in $\mathcal{L}$ by a similar argument. This shows the claim.
Since $(\overline{f},\phi)$ and $(\overline{g},\psi)$ are pre-morphisms, the following diagrams commute for each $k>n$:
$V_{\phi(k)}$$W_{k}$$V_{\phi(n)}$$W_{n}$$\scriptstyle{f_{k}}$$\scriptstyle{\widehat{p}_{\phi(k),\phi(n)}}$$\scriptstyle{\widehat{q}_{k,n}}$$\scriptstyle{f_{n}}$$W_{\psi(k)}$$V_{k}$$W_{\psi(n)}$$V_{n}$$\scriptstyle{g_{k}}$$\scriptstyle{\widehat{q}_{\psi(k),\psi(n)}}$$\scriptstyle{\widehat{p}_{k,n}}$$\scriptstyle{g_{n}}$
Let $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$ be defined by $r_{n}=F(f_{n})$ and $s_{n}=F(g_{n})$ where $F$ is the functor from Lemma 4.6. By that lemma we obtain commutative diagrams of surjections between sets
$T_{\phi(k)}$$U_{k}$$T_{\phi(n)}$$U_{n}$$\scriptstyle{r_{k}}$$\scriptstyle{p_{\phi(k),\phi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{r_{n}}$$U_{\psi(k)}$$T_{k}$$U_{\psi(n)}$$T_{n}$$\scriptstyle{s_{k}}$$\scriptstyle{q_{\psi(k),\psi(n)}}$$\scriptstyle{q_{k,n}}$$\scriptstyle{s_{n}}$
Let $\Phi\colon[T]\to[U]$ and $\Psi\colon[U]\to[T]$ be defined by
$\Phi(x)=\bigcup_{n}r_{n}(x\upharpoonright\phi(n))\qquad\Psi(y)=\bigcup_{n}s_{n}(y\upharpoonright\psi(n)).$
The $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$4.2) and 4.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎